OpenStudy (unklerhaukus):

\[(\forall n\in \mathbb Z)[n|n^2]\]

OpenStudy (unklerhaukus):

t/f

Parth (parthkohli):

What does "|" mean in logic?

OpenStudy (anonymous):

divides

Parth (parthkohli):

The division thing? Yes.

Parth (parthkohli):

Except for 0.

OpenStudy (swissgirl):

True

Parth (parthkohli):

What if \(\rm n = 0\)?

OpenStudy (swissgirl):

n^2=n*n so (n*n)/n=n

Parth (parthkohli):

False.

OpenStudy (swissgirl):

Truesomeness

Parth (parthkohli):

Though the statement is true for \(\mathbb{Z}^+\)

OpenStudy (anonymous):

Dont read :\[n\mid n^2\]as division. By definition, \[a\mid b \iff b=ak ,k\in \mathbb{Z}\]there is no division taking place. By this definition, 0 does divide 0^2 since 0^2=0*0

Parth (parthkohli):

I've always heard, zero cannot divide anything.

OpenStudy (anonymous):

You are thinking of fractions. Notice the definition doesnt contain anything about fractions.

OpenStudy (anonymous):

its a statement only about multiplication.

Parth (parthkohli):

Oh.

OpenStudy (unklerhaukus):

\[a|b \iff \exists q[b=aq],a\neq 0\]

OpenStudy (anonymous):

ah, then i am mistaken. if a cant be zero then what i posted is wrong.

Parth (parthkohli):

So, in fact, we do have fractions in the definition.

OpenStudy (anonymous):

hmmm, i still wouldnt say there are fractions. This is generally the way they talk about division in Rings, where only multiplication and addition are defined. But yes, a cannot be zero. http://primes.utm.edu/glossary/xpage/divides.html

OpenStudy (unklerhaukus):

\(a|b\) denotes \(b\) is divisible by \( a \)